1. Gibson Manufacturing Corporation expects to sell the following number of units of steel cables at the prices indicated, under three different scenarios in the economy. The probability of each outcome is indicated.

Outcome Probability Units Price
A 0.70    280     $31  
B 0.20    480     46
C 0.10    730     56

What is the expected value of the total sales projection?

Total expected value           $   

2. Sales for Ross Pro’s Sports Equipment are expected to be 41,000 units for October. The company likes to maintain 10 percent of unit sales for each month in ending inventory (that is, end of October). Beginning inventory for October is 9,000 units.

How many units should the firm produce for the coming month?

Units to be produced -

3. Garza Electronics expects to sell 700 units in January, 450 units in February, and 2,000 units in March. January’s beginning inventory is 2,000 units. Expected sales for the whole year are 19,200 units. Garza has decided on a level monthly production schedule of 1,600 units (19,200 units/12 months = 1,600 units per month). What is the expected end-of-month inventory for January, February, and March?

Ending
inventory
  January    
  February    
  March    
Answer #1
1. What is the expected value of the total sales projection?
Total expected value = (0.70 x 280 x $31) + (0.20 x 480 x $46) + (0.10 x 730 x $56) = $14,580
2. How many units should the firm produce for the coming month?
Units to be produced = 41,000 + (41,000 x 10%) - 9,000 = 36,100
3. What is the expected end-of-month inventory for January, February, and March?
Ending inventory
January 2,000 + 1,600 - 700 = 2,900
February 2,900 + 1,600 - 450 = 4,050
March 4,050 + 1,600 - 2,000 = 3,650